# pasta: Pattern Analysis for Spatial Omics Data

### v2

@@ Results
Many imaging platforms measure patterns in a field of view (FOV) and collect several such FOVs per sample. In some technologies, there are no gaps between the individual FOVs and they can be assembled into one big image. If there are gaps, consecutive FOVs can either be stitched together (which can be computationally complex) or analysed individually [aspSpatiallyResolvedTranscriptomes2020, mohammadiFastRobustFeaturebased2024]. When analysing individual FOVs, we do not observe the entire pattern but rather a subset of the bigger pattern. This is called window sampling because the window is just a sample of the bigger point pattern. A related concept is the small world model, which describes that points can only be observed in a finite world and not beyond these boundaries [baddeleySpatialPointPatterns2015, pp. 143ff.]. Often, the distinction between the window sampling and small world model concepts is not clear. An example is the arrangement of the epithelium, which can can be imaged with several FOVs (window sampling) but there are no cells expected outside of the epithelial boundary (small world) [garrido-trigoMacrophageNeutrophilHeterogeneity2023]. Therefore, in cell biology, we often encounter a mixture of these two concepts.
Apart from the window of measurement, we need to make assumptions on the statistical properties of the point pattern, including, most importantly, whether points can be considered homogeneous or not. Homogeneity (see Box Assumptions) assumes that the number of points in a given region is proportional to its area , for arbitrary ; i.e., homogeneity refers to a uniform intensity of points across the window of measurement. If that is not the case, the point process is said to be inhomogeneous [baddeleySpatialPointPatterns2015, p.132-133]. This difference has important implications for the analysis of a process. For example, when adjusting for varying local intensities, the process could be found to be inhomogeneous rather than genuinely clustered. This is called the confounding between intensity and interaction (see box Assumptions) [baddeleySpatialPointPatterns2015, pp.151-152].
Options for analysis
Point pattern analysis offers a broad spectrum of options. Most often, we compare patterns in 2D (e.g., a single histological slice), but the methods are applicable with some adjustments to 3D patterns (e.g., confocal microscopy volumes) [baddeleySpatialPointPatterns2015, p. 651], with some increased computational cost [baddeleyAnalysisThreeDimensionalPoint1993].
As mentioned above, points can carry marks; the nature of these marks determines what options are available to the analyst. For example, many point pattern methods are designed for categorical data, such as cell types or spatial domain labels [baddeleySpatialPointPatterns2015, p.147] and can be adapted to the analysis of continuous marks [baddeleySpatialPointPatterns2015, pp.644 ff.]. Model-based approaches for spatial data might also be of interest. This is an entire subfield on its own with many options available [baddeleySpatialPointPatterns2015, pp.297 ff.][pebesmaSpatialDataScience2023].
When comparing categorical marks, the analyst must decide whether to compare a single type (within-type comparison) or between a pair of types (cross-type comparison). For example, a within-type comparison could use the pattern of epithelial cells in a microscopy image; the correlation, spacing or distances between epithelial cells could be summarised. In contrast, a cross-type analysis could consider the correlation between epithelial cells and B cells [baddeleySpatialPointPatterns2015, pp. 561ff.].
Examples
A well-known point pattern summary is Ripley’s -function, defined as the expected number of points falling in a specified radius around an arbitrary point, averaged over all points in the observation window (e.g., FOV) (potentially adjusted for edge effects; see box Assumptions). Ripley’s -function thus quantifies the correlation structure of a point pattern, with within-type and cross-type versions. A point pattern where points have no preference for location shows complete spatial randomness (CSR). This means that the number of points that fall in an arbitrary region follow a Poisson distribution with a rate parameter that is proportional to the size of the region. A clustered process would have values of the -function larger than the Poisson process at that specific scale; values of smaller than the Poisson process indicate a self-inhibiting process (i.e., where points repel each other) [ripleySecondOrderAnalysisStationary1976, ripleyModellingSpatialPatterns1977][baddeleySpatialPointPatterns2015, pp.203 ff.].
Ripley’s -function can be variance stabilised by taking the square root, known as Besag’s -function [besag1977contribution]. It is a global estimate that measures correlation across an entire observation window and can be used instead of the standard -function [baddeleySpatialPointPatterns2015, p.402], [baddeleySpatstatPackageAnalyzing2005].
Figure 2A-C show point patterns of mature oligodendrocytes over three different serial sections of the same mouse hypothalamic preoptic region (the numbers , , indicate the -axis position of the slices: one slice every 50 m along the anterior-posterior axis [moffittMolecularSpatialFunctional2018]). Figure 2D shows the homogeneous global -function, which quantifies the correlation within each slice and the respective -function for a Poisson process (dashed line). The homogeneous -function indicates strong clustering, even for the slice, however, the inhomogeneous -function, which makes a local adjustment for intensity, indicates a level of homogeneity very similar to a (completely spatially random) Poisson process (Supp. Figure S1A-C vs. D-F). This is an example of intensity vs. interaction confounding (see box Assumptions). If slice should therefore be interpreted as clustered, depends on the assumption of homogeneity for that slice. In addition, there are differences between the inhomogeneity corrections in Figure S3; the corrections are discussed in more detail in the vignette.
Local versions of the various summary functions exist. The LISA (local indicators of spatial association) framework provides local versions for many spatial statistics (including point pattern and lattice data analysis) [anselinLocalIndicatorsSpatial1995]. For example, Figure 2E shows LISA curves for the slice [getisSecondOrderNeighborhoodAnalysis1987, anselinLocalIndicatorsSpatial1995]. That is, for each point, we calculate a local -function. These curves can be analysed with functional principal component analysis (fPCA), which allows to extract the main functional modes of variation [ramsayPrincipalComponentsAnalysis2005]. The idea to perform fPCA on spatial statistics summaries has been used in ecological and spatial omics data [illianPrincipalComponentAnalysis2006, vuFunSpaceFunctionalSpatial2023, vuSPFSpatialFunctional2022]. Figure 2F shows the scores of the first two functional PCs. There are two main clusters, representing either the physically clustered oligodendrocytes or those in low-density regions.
Challenges and limitations
Analysing biological samples formally as point patterns comes with several challenges and limitations. The first and most important is viewing the biological sample as the realisation of a point process. It is generally not easy to define whether the entire pattern of several cell types should be represented as a single point process (multitype analysis). Analysing the cell types individually and considering them as separate point processes might underestimate dependencies among the cells (multivariable analysis) [baddeleySpatialPointPatterns2015, pp. 565 ff.]. If, as in our example in Figure 2, the sections are far apart from each other, it can be justified to analyse the processes in isolation as a multivariable analysis. However, if we consider cells in two adjacent slices, these have probably been formed by one biological process and should not be analysed in isolation.
Obtaining a representative sample of a point pattern is also a challenge. In spatial omics, we are often provided with FOVs that an experimentalist has selected from a much larger region or that stem from the technology itself. Such a sample might have different characteristics compared to the point pattern of the larger region, as FOVs are sometimes selected based on morphological properties (e.g., H&E staining). In cases where we have a small world scenario, we would want to limit our observation window to the full region where the points can occur.
Lattice Data Analysis
Definitions
In contrast to point pattern-based methods, analysts could take the observation-based view and regard the location of cells or spots as a fixed (regular or irregular) lattice and consider the corresponding marker expression at each cell or spot. Several HTS-based spatial transcriptomics technologies produce data on a regular lattice (i.e., evenly spaced spots or beads of uniform size and shape), whereas imaging-based technologies yield irregular lattice structures (i.e., segmentation resulting in variable cell sizes and shapes, and non-uniform spacing) [zuurAnalysingEcologicalData2007, pp. 321ff.] [anselin2024introduction].
Assumptions
Once we have recorded the outline and arrangement of the spatial units in the lattice, we can specify the strength of spatial relationships between each unit. Each pair of units is assigned a weight: the stronger the connection between units, the higher the weight. The collection of all weights between locations and , , forms the weight matrix (see box Options for analysis). The construction of the weight matrix is critical for all downstream analyses as it encodes the spatial relationship between the units in the lattice [zuurAnalysingEcologicalData2007, pp. 321ff.] and there exist various strategies: contiguity-based (i.e., in direct contact), graph-based, distance-based, or higher-order neighbours [getisSpatialWeightsMatrices2009, zuurAnalysingEcologicalData2007, pebesmaSpatialDataScience2023].
Options for analysis
A common exploratory analysis for lattice data is based on the concept of spatial autocorrelation. Spatial autocorrelation measures the degree of association of features, e.g., the expression of a gene, that are assumed to be dependent in space. For each combination of spatial locations, a measure of association is calculated and scaled by the weight of the connection [anselin2024introduction]. This approach considers both the proximity (via the weight matrix) and the characteristics of the locations (via the metric) [zuurAnalysingEcologicalData2007, pp. 327ff.] [pebesmaSpatialDataScience2023, pp. 209ff.] [getisAnalysisSpatialAssociation1992]. Over the years, various association measures have emerged, each specific to an aspect of the data. While many allow for univariate comparisons, in which one variable (e.g., expression of a gene) is compared over multiple locations, others allow for the comparison of two (bivariate) or more (multivariate) features at once [pebesmaSpatialDataScience2023, mosesVoyagerExploratorySinglecell2023]. For most metrics, the features are assumed to be continuous measurements. However, metrics allow for the measurement of spatial autocorrelation of categorical variables (e.g., join count statistics). In addition to the number of features and the type of measurement that can be compared in the association score, there is the to consider: global spatial autocorrelation metrics estimate the average level of spatial autocorrelation across all locations whereas local measures give a statistic at each location. The global statistic can be seen as the weighted sum of its respective local statistics [anselinLocalIndicatorsSpatial1995].
Example
Moran’s is a prominent example of a spatial autocorrelation metric. It can be interpreted as the mean Pearson correlation between values at locations and the mean value of the corresponding neighbours [moranNotesContinuousStochastic1950]. A positive value of Moran’s indicates that locations with similar values are clustered while a negative value indicates the clustering of dissimilar values, giving a measure of spatial heterogeneity. The global value of Moran’s is bounded by and with an expected value under spatial randomness near for large , (). Figure 3A shows evidence for spatial autocorrelation of Nrgn expression in mouse brain tissue; the (global) Moran’s value is (p-value essentially zero). Global measures can be used to identify spatially variable genes when compared to the global autocorrelation values of different genes [weberNnSVGScalableIdentification2023, chenEvaluatingSpatiallyVariable2024]. Furthermore, there is a local version of Moran’s (interpretation is analogous to the global version). Figure 3B shows the estimate of local Moran’s at each measured spot (n.b., the local version is not bounded by and ). Figure 3C shows the corresponding adjusted p-values (testing for non-zero spatial autocorrelation). When interpreting local autocorrelation measures, it is important to consider both the effect size estimates and the significance level. Since the significance level is calculated for each spot separately, it is recommended to adjust for multiple testing (e.g., Benjamini and Hochberg [benjamini_hochberg1995fdr]). Overall, local Moran’s statistics reveal locations that have similar values to their neighbours. However, as can be seen in Figure 3A-C regions with the highest and most significant local Moran’s value lie in the part of the tissue where expression of gene Nrgn is very low amongst neighbours. Notably, the local Moran’s measure is both dependent on the value of the log-transformed counts and the similarity among neighbours (c.f., Supplementary Figure S2A-B). For the analysis of categorical data, there are metrics such as the join count statistics [daleSpatialAnalysisGuide2014, pp. 141ff.]. In essence, the join count statistic calculates the frequency of categories among neighbours and compares this value with a theoretical distribution or permutations of the labels to get a significance score. This can be used to investigate the interactions between cell types in a lattice since the distance can be specified when constructing the neighbourhood graph.
Challenges and limitations
The choice of the weight matrix poses a challenge in the analysis of lattice data, as it might influence downstream analyses and conclusions. A common choice is to give a non-zero weight only to direct neighbours of a spatial unit (i.e., contiguity-based). Neighbours based on graphs or on distance [pebesmaSpatialDataScience2023, pp. 191 ff.] are also used; the open question is which is best suited for spatial omics data. If one uses contiguity-based neighbours in imaging-based spatial transcriptomics, results ultimately depend heavily on the accuracy of the cell segmentation, which is known to be challenging [stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023, wangCellSegmentationImage2019]. If neighbours beyond the adjacent ones are used, the scale (e.g., at which cells interact still) needs to be specified. Figure 3C-D show the difference between the local Moran’s calculation when based on contiguity-based neighbours in D, the 10 nearest neighbours in E or neighbours within a 1000 pixel distance () in F. While the overall differences are small, some cells show different local Moran’s values. Smoothing of the local Moran’s values occurs when more neighbours are considered (Supplementary Figure S2C). Supplementary Figure S2A and B further show that for some cells, no contiguous neighbours were found, which results in a zero estimate of the local Moran’s . Since cells do not function in isolation but form complex anatomical structures including extracellular components, an open question is whether analyses improve when weight matrix construction takes anatomical structures or regions into account. Overall, it remains to be investigated how much the construction of the weight matrix influences downstream analyses in spatial omics data.
Short guide to spatial analysis of biological data
The large variety of options for analysis in spatial statistics can be challenging to navigate. A few fundamental steps should help selecting appropriate exploratory analysis. First, analysts should consider their technology and what options they have (i.e., imaging-based or HTS-based approaches). As seen in Figure 1, the data representation leads to either point pattern or lattice data analysis. If both streams are possible, the decision depends on the research question. In point pattern analysis, the process leading to the locations of individual points is the primary interest. It allows us to make statements such as whether a point process exhibits clustering or if two cell types tend to interact more than randomly expected. In lattice data analysis, however, we are not directly interested in the arrangement of points, but rather locations are important to define the spatial relationship between locations (i.e., via the weight matrix). We study the interaction of features, such as gene expression, given their spatial location. For example, if ligand-receptor dynamics are of interest, the weight matrix can be specified in a way that captures spatial associations of gene expression between direct neighbours or at a specified scale.
@@
Second, the observation scale has to be chosen. Sometimes, this is a technological scale such as multiple FOVs in imaging. Figure 4A shows a single FOV with islets from the human pancreas acquired using IMC [damondMapHumanType2019]. An analysis considering the entire FOV will tell us about the processes that gave rise to the spatial distribution of the endocrine pancreas (islets). As we see in Figure 4B, a Ripley’s -function under a homogeneity assumption indicates that in this observation window, endocrine cells are clustered within pancreatic islets. However, there might also be a relevant biological scale where we subset our observation window to the islets themselves (Figure 4D). In this case, Ripley’s -function indicates a homogeneous distribution of the islets cells within the observation window (Figure 4E). Overall, the correct scale needs to be chosen in accordance with the research question, since the interpretation depends on it.
In point pattern analysis, the analyst needs to decide whether (and how) to control for inhomogeneous intensities, leading back to potential confounding of intensity and interaction [baddeleySpatialPointPatterns2015, pp. 151 ff.]. Analysts can look at smoothed density estimates to inspect potential inhomogeneity prior to analysing the pattern more formally. Furthermore, it depends again on the research question and biological domain knowledge whether the assumption of homogeneity is appropriate. While tissues such as the adipose tissue form regular structures that could be assumed to be homogeneous, other tissues such as stratified epithelial structures appear to be more inhomogeneous. As shown in Figure 4, the homogeneous and inhomogeneous variants of the spatial summaries can lead to different interpretations.
@@
A fundamental choice in lattice data analysis is the construction of the weight matrix. As seen in Figure 3A-B, there are differences in the interpretation of results between different weight matrix choices. The appropriate weight matrix design depends on the research question. For instance, if the feature of interest is a ligand that requires direct cell contact, using contiguous neighbours is a reasonable choice. If the feature of interest is a molecule that diffuses to nearby cells (e.g., paracrine signalling), a nearest neighbour or distance-based approach should be considered for constructing the neighbourhood matrix.

### v3

@@ Introduction
Many imaging platforms measure patterns in a field of view (FOV) and collect several such FOVs per sample. In some technologies, there are no gaps between the individual FOVs so they can be assembled into one image. If there are gaps, consecutive FOVs can either be stitched together (which can be computationally complex) or analyzed individually [aspSpatiallyResolvedTranscriptomes2020, mohammadiFastRobustFeaturebased2024]. When analyzing individual FOVs, not the entire pattern but rather a subset of the full pattern is observed: this is so-called window sampling because the window is a sample of the full point pattern. A related concept is the small world model, which describes that points can only be observed in a finite world and not beyond these boundaries [baddeleySpatialPointPatterns2015, pp. 143ff., p. 560]. Often, the distinction between the window sampling and small world model concepts is not clear in histological sections. An example is the arrangement of the epithelium, which can can be imaged with several FOVs (window sampling) but there are no cells expected outside of the epithelial boundary (small world) [garrido-trigoMacrophageNeutrophilHeterogeneity2023]. Therefore, in spatial omics, we often encounter a mixture of these two scenarios.Obtaining a representative sample of a point pattern is also a challenge. In spatial omics, we are often provided with FOVs that an experimentalist has selected from a much larger region or that are the result of the data acquisition (e.g., FOVs in CosMx [heHighplexImagingRNA2022]). Such a sample might have different characteristics compared to the point pattern of the larger region, as FOVs are sometimes selected based on morphological properties (e.g., H&E staining). In cases where we have a small world scenario, we would want to limit our observation window to the full region where the points can occur (c.f., Supplementary Figure S3).
Importance of scale and uniformity in point pattern analysis
@@
Apart from the window of measurement, we need to make assumptions on the statistical properties of the point pattern, including, most importantly, whether points can be considered homogeneous or not. Homogeneity (see Box Importance of scale and uniformity in point pattern analysis) assumes that the number of points in a given region is proportional to its area, i.e., homogeneity refers to a uniform intensity of points across the window of measurement. Otherwise, the point process is inhomogeneous [baddeleySpatialPointPatterns2015, p.132-133]. This difference has important implications for the analysis of a point process. For example, when adjusting for varying local intensities, a process could be interpreted as inhomogeneous rather than genuinely clustered. This is called the confounding between intensity and interaction (see Box Importance of scale and uniformity in point pattern analysis) [baddeleySpatialPointPatterns2015, pp.151-152]. Figure 4A shows a single FOV with islets from the human pancreas acquired using IMC [damondMapHumanType2019]. An analysis considering the entire FOV will tell us about the processes that gave rise to the spatial distribution of the endocrine pancreas (islets). As we see in Figure 4B, a Ripley’s -function under a homogeneity assumption indicates that in this observation window, endocrine cells are clustered within pancreatic islets (the line lies above the theoretical CSR line). Figure 4B corrects for inhomogeneity of islets cells in the total fov and indicates CSR at . Note that the calculation of the inhomogeneous version varies with the estimation of the intensity function and the border correction (c.f., Supplementary Figure S5), especially when the data is not correlation-stationary [baddeleySpatialPointPatterns2015, pp. 321ff.]. However, there might also be a relevant biological scale where we subset our observation window to the islets themselves (Figure 4D). In this case, Ripley’s -function indicates a homogeneous distribution of the islet cells within the observation window (Figure 4E, the line follows the theoretical CSR line). Overall, the correct scale needs to be chosen in accordance with the research question, since the interpretation depends on it.
Regular and irregular lattices
Several HTS-based spatial transcriptomics technologies produce data on a regular lattice that gives evenly-spaced spots or beads of uniform size and shape, whereas imaging-based technologies yield irregular lattice structures with variable cell sizes and shapes, and non-uniform spacing [zuurAnalysingEcologicalData2007, pp. 321ff.] [anselin2024introduction]. Once we have recorded the outline and arrangement of the spatial units in the lattice, we can specify the strength of spatial relationships between each unit. Each pair of units is assigned a weight: the stronger the connection between units, the higher the weight. The collection of all weights between locations forms the weight matrix (see Box Definition of the neighbourhood is critical in lattice data analysis).
Spatial autocorrelation is a measure of similarity
A common exploratory analysis for lattice data is based on the concept of spatial autocorrelation. Spatial autocorrelation measures the degree of association of features, e.g., the expression of a gene, that are assumed to be dependent in space. For each combination of spatial locations, a measure of association is calculated and scaled by the weight of the connection [anselin2024introduction]. This approach considers both the proximity (via the weight matrix) and the characteristics of the locations (via the metric) [zuurAnalysingEcologicalData2007, pp. 327ff.] [pebesmaSpatialDataScience2023, pp. 209ff.] [getisAnalysisSpatialAssociation1992]. When interpreting (local) autocorrelation measures, it is important to consider both the effect size estimates and the significance level. Since the significance level is calculated for each spot separately, it is recommended to adjust for multiple testing (e.g., Benjamini and Hochberg [benjamini_hochberg1995fdr]). Overall, local Moran’s statistics reveal locations that have similar values to their neighbours. However, in mouse brain, regions with the highest and most significant local Moran’s value lie where expression of gene Nrgn is very low among neighbours (Figure 5A-B). Notably, the local Moran’s measure is both dependent on the value of the log-transformed counts and the similarity among neighbours (c.f., Supplementary Figure S4A-B).Metrics such as the Moran’s scatter plot can be useful to classify the type of interaction [anselin2019moran]. The Moran’s scatter plot compares the values of each spot with the average of it’s neighbours as defined in the weight matrix. This information can be used to classify each point into respective categories (c.f., Supplementary Figure S4E)). For example, in Figure 5 C the clusters denoted “high-high” refer to regions where high expression values are surrounded by high values, while “low-low” clusters indicate that low values are surrounded by low values (for which the local Moran’s value can be high as well). “High-low” and “low-high” values denote heterogeneous regions.
Definition of the neighbourhood is critical in lattice data analysis
The construction of the weight matrix is critical for all downstream analyses as it encodes the spatial relationship between the units in the lattice [zuurAnalysingEcologicalData2007, pp. 321ff.]. Several construction strategies exist: contiguity-based (i.e., in direct contact), graph-based, distance-based, or higher-order neighbours [getisSpatialWeightsMatrices2009, zuurAnalysingEcologicalData2007, pebesmaSpatialDataScience2023]. Neighbours based on graph adjacency or on distance [pebesmaSpatialDataScience2023, pp. 191 ff.] are also used; the open question is which is best suited for spatial omics data. If one uses contiguity-based neighbours in imaging-based spatial transcriptomics, results ultimately depend heavily on the accuracy of the cell segmentation, which is known to be challenging [stringerCellposeGeneralistAlgorithm2021, greenwaldWholecellSegmentationTissue2022, petukhovCellSegmentationImagingbased2022, architSegmentAnythingMicroscopy2023, wangCellSegmentationImage2019]. If neighbours beyond the adjacent ones are used, the scale (e.g., at which cells interact) needs to be specified. Figure 5C-D show the difference between the local Moran’s calculation when based on contiguity-based neighbours in D, the 10 nearest neighbours in E or neighbours within a 1000 pixel distance ( ) in F. While the overall differences are small, some cells show different local Moran’s values. Smoothing of the local Moran’s values occurs when more neighbours are considered (Supplementary Figure S4C). Furthermore, Supplementary Figure S4A and B show that for some cells, no contiguous neighbours were found, which results in a zero estimate of the local Moran’s . Cells do not function in isolation but form complex anatomical structures including extracellular components. Whether analyses improve when weight matrix construction takes such anatomical structures or regions into account is to be determined. Furthermore, it remains to be investigated how much the construction of the weight matrix influences downstream analyses in spatial omics data.
pasta highlights various spatial statistics analyses with inline code in R and Python
The collection of vignettes on our webpage (https://robinsonlabuzh.github.io/pasta) showcases the analysis of data from multiple spatial omics technologies, where concepts and assumptions are discussed in detail, with inline R and Python code. The design of the webpage follows the outline of Figure 1. First, we consider the technology and categorize into a HTS-based or imaging-based approach. Second, we distinguish between univariate (one type of mark) and multivariate (more than one type of mark) methods and outline different methods in specific vignettes with an in-depth discussion of the methods, e.g. providing mathematical definitions and formulae. In addition, we provide overview vignettes summarizing the two approaches.
To ensure usability, we showcase exploratory spatial statistics methods integrated into existing bioinformatics ecosystems. In Python, we use anndata objects from scanpy, allowing direct usage with SpatialData as well [virshup2024anndata, wolf2018scanpy, marconato2025spatialdata]. In R, we integrate with SpatialExperiment and SpatialFeatureExperiment objects from Bioconductor [righelliSpatialExperimentInfrastructureSpatiallyresolved2022, mosesVoyagerExploratorySinglecell2023, gentleman2004bioconductor]. The packages Voyager and spatialFDA provide wrappers to compute both lattice and point pattern metrics on Bioconductor objects [mosesVoyagerExploratorySinglecell2023, emonsSpatialFDAToolSpatial2025]. Furthermore, our vignette offers examples how to convert between anndata objects in Python and SpatialExperiment, Seurat objects objects in R [righelliSpatialExperimentInfrastructureSpatiallyresolved2022, haoDictionaryLearningIntegrative2023]. Using these conversions, we are enabling analysis for users of both R and Python.
Overall, our resource Pattern Analysis for SpaTial omics datA (pasta) highlights the usefulness and transferability of existing exploratory spatial statistics approaches in the context of spatial tissue profiling.
Spatial statistics is used in existing spatial omics data analysis tools
There are already several tools for omics analysis that employ spatial statistics. Among others, Voyager is a framework and a collection of use cases for lattice data written in both R and Python [mosesVoyagerExploratorySinglecell2023]; it provides data structures and functionalities to compute spatial statistics in molecular biology. Voyager comes with comprehensive vignettes for several spatial profiling technologies. The methodological focus of their vignettes is on lattice data analysis and the spdep geospatial package [bivandPackagesAnalyzingSpatial2022]. Since Voyager offers efficient implementations for lattice data analysis in R, we build on their framework in our vignettes for the lattice data component. Other methods written in R include: SPIAT, which contains various tools for spatial analysis, such as: neighbourhood analysis, local metrics and heterogeneity scores etc. [fengSpatialAnalysisSPIAT2023]; scFeatures, a toolbox comprising lattice data analysis functionality such as Moran’s and point pattern metrics as the -function [caoScFeaturesMultiviewRepresentations2022]; spicyR, which uses a compressed version of the global -function for cross-sample comparison [caneteSpicyRSpatialAnalysis2022]; lisaClust, a spatial domain detection method via LISA (local indicators of spatial association) curves [patrickSpatialAnalysisHighly2023]; mxfda, that uses point pattern summary functions for survival analysis [wrobelMxfdaComprehensiveToolkit2024]; spatialDM uses global and local bivariate Moran’s to score co-localization of ligand-receptor pairs [liSpatialDMRapidIdentification2023]; and MERINGUE, which uses nearest-neighbour autocorrelation for spatially variable gene selection [millerCharacterizingSpatialGene2021]. Methods in Python include squidpy, a package with various spatial statistics tools for both point pattern and lattice data [pallaSquidpyScalableFramework2022]; PySAL a general Python library for spatial analysis [reyPySALPythonLibrary2010]; Bento a toolkit for analysis of subcellular features using different spatial statistics [mahBentoToolkitSubcellular2024]; SpottedPy a package designed to identify tumor hotspots using spatial statistics [withnellSpottedPyQuantifiesRelationships2024]; MuSpAn a toolbox for multiscale spatial analysis using spatial statistics [bullMuSpAnToolboxMultiscale2024] and spatiomic a Python library for the analysis of pixel based spatial data using spatial correlation measures [kuehlSpatiomicScalableSpatial2025].
For a detailed view on general methods development in the field of spatial transcriptomics, we refer readers to the online Supplementary table of the museum of spatial transcriptomics paper [mosesMuseumSpatialTranscriptomics2022].
Short guide to spatial analysis of spatial omics data
Step 1 – The data representation
First, analysts should consider their technology and what options they have (i.e., imaging-based or HTS-based approaches). As seen in Figure 1, the data representation leads to either point pattern or lattice data analysis. If both streams are possible, the decision depends on the research question. In point pattern analysis, the process leading to the locations of individual points is the primary interest. It allows us to make statements such as whether a point process exhibits clustering or if two cell types tend to interact more than expected under CSR. In lattice data analysis, however, we are not directly interested in the arrangement of points, but rather locations are used to define the structure of the neighbourhood (i.e., via the weight matrix). We study the interaction of features, such as gene expression levels, given their spatial location.
Step 2 – Choice of feature and nature of the spatial interaction
Second, the feature has to be considered. Most often this will be either gene or protein expression (continuous feature) or cell type or other labels (categorical feature). As seen in Figure 2, lattice data analysis offers many more methods for continuous features, whereas the focus of point pattern analysis is categorical features. Next, the scientific question guides the choice of the analysis. For example, if we are interested in the expression of one gene in space, this is a univariate continuous comparison. Another example would be the comparison of the co-localization of two cell types, which is a bivariate categorical comparison.
@@
A fundamental choice in lattice data analysis is the construction of the weight matrix. As seen in Figure 5D-F, there are differences in the interpretation of results between different weight matrix choices. The appropriate weight matrix design depends on the research question. For instance, if the feature of interest is a ligand that requires direct cell contact, using contiguous neighbours is a reasonable choice. If the feature of interest is a molecule that diffuses to nearby cells (e.g., paracrine signalling), a nearest neighbour or distance-based approach should be considered for constructing the neighbourhood matrix.
Figure 1 helps with the navigation of both data modalities and analysis options.
Step 3 – Observation scale
Third, the observation scale has to be chosen. Sometimes, a scale given by a technology such as a FOV might be appropriate. In other cases, the observation scale should be given by the biological question. Recall the example of Figure 4 where the scale effectively determines whether the point pattern of islet cells is considered to be homogeneous or clustered.
In lattice data analysis the scale is directly connected to the weight matrix. For example, if ligand-receptor co-localizations are of interest, the weight matrix can be specified in a way that captures spatial associations of gene expression between direct neighbours or at a specified scale. Generally, the concept of spatial autocorrelation is based on Tobler’s famous first law of geography, which states that “everything is related to everything else, but near things are more related than distant things” [toblerComputerMovieSimulating1970]. This statement is not generally applicable in biology where we face highly compartmentalized anatomical structures. In analysis, it might therefore be important to account for this compartmentalization and define the neighbourhood accordingly or perform analysis that is aware of different compartments.
Step 4 – Technical considerations
In point pattern analysis, the analyst needs to decide whether (and how) to control for inhomogeneous intensities, leading back to potential confounding of intensity and interaction [baddeleySpatialPointPatterns2015, pp. 151 ff.]. Analysts can look at smoothed density estimates to inspect potential inhomogeneity prior to analyzing the pattern more formally. Furthermore, it depends again on the research question and biological domain knowledge whether the assumption of homogeneity is appropriate. While some tissues form regular structures that could be assumed to be homogeneous, other tissues appear to be more inhomogeneous. As shown in Figure 4, the homogeneous and inhomogeneous variants of the spatial summaries can lead to different interpretations.